11. Find the number of ways that four flags can be displayed on a flag pole, one above the other, if ten different flags are available.
Answer by stanbon(48522)

Find the number of ways that four flags can be displayed on a flag pole, one above the other, if ten different flags are available.
---------------------------
Order makes a difference:
10P4 = 10!/(10-4)! = 10*9*8*7 = 5040
========================================
Cheers,
Stan H.